### v1

# Rethinking the Relationship between the Power Law and Hierarchical Structures

Abstract
Statistical analysis of corpora provides an approach to quantitatively investigate natural languages. This approach has revealed that several power laws consistently emerge across different corpora and languages, suggesting the universal principles underlying languages. Particularly, the power-law decay of correlation has been interpreted as evidence for underlying hierarchical structures in syntax, semantics, and discourse. This perspective has also been extended to child languages and animal signals. However, the argument supporting this interpretation has not been empirically tested. To address this problem, this study examines the validity of the argument for syntactic structures. Specifically, we test whether the statistical properties of parse trees align with the implicit assumptions in the argument. Using English corpora, we analyze the mutual information, deviations from probabilistic context-free grammars (PCFGs), and other properties in parse trees, as well as in the PCFG that approximates these trees. Our results indicate that the assumptions do not hold for syntactic structures and that it is difficult to apply the proposed argument to child languages and animal signals, highlighting the need to reconsider the relationship between the power law and hierarchical structures.
11footnotetext: This work was conducted while KN and KK were at the University of Tokyo.
1 Introduction
Natural languages exhibit characteristic statistical properties. Quantifying these properties through corpus-based analyses is crucial for understanding natural languages, providing insights into the mechanisms that give rise to the properties. Previous research has extensively explored this approach, uncovering various power laws, such as Zipf’s law (Zipf, 2016) and Heaps’ law (Heaps, 1978), that consistently appear across different corpora and languages. These findings suggest the existence of universal principles of languages.
This study focuses on the power-law decay of correlation. In languages, distant elements can remain strongly correlated. For example, subject–verb agreement in English can occur at arbitrarily large distances, and a single sentence can have semantic influence over an entire text. Previous studies have measured correlations between words, characters, phones, and other elements using metrics such as correlation functions and mutual information (MI). These studies have consistently found that the correlation decays according to a power-law function of the distance between elements (Li, 1989a; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023; Nakaishi et al., 2024). This phenomenon indicates that the correlation is significant even at large distances.
Moreover, the power-law decay of correlation has been interpreted as evidence supporting the linguistic view that languages are organized hierarchically in syntax, semantics, and discourse (Alvarez-Lacalle et al., 2006; Altmann et al., 2012; Lin and Tegmark, 2017). This perspective has also been extended beyond adult human languages. Previous research has observed similar power-law behaviors in child languages and in animal signals, arguing that these may also possess underlying hierarchical structures (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
This interpretation is justified by the argument of Lin and Tegmark (2017). They showed that the decay of correlation follows a power law in sequences generated by probabilistic context-free grammars (PCFGs; Jelinek et al., 1992), which are a simple model for generating hierarchical tree structures. To be more precise, they proved that in PCFGs, the correlation decays exponentially in structures and assumed that the sequential distance grows exponentially with the structural distance. These give rise to the power-law decay of correlation in sequences.
However, this argument rests on several implicit assumptions. Previous studies have not tested their validity. To address this, we examine whether these assumptions hold for hierarchical structures in syntax, i.e., parse trees. Although parsing schemes vary in detail, syntactic structures have been well formalized, resulting in the availability of large-scale treebanks. This enables us to perform statistical analyses on treebanks to test these assumptions.
Specifically, we formalize three key questions essential for examining the validity of the assumptions: (i) Does the correlation decay exponentially in structures?; (ii) Does the sequential distance grow exponentially with the structural distance?; (iii) Are PCFGs adequate models for the hierarchical structures of natural languages? To answer these questions, we analyze treebanks.
As a result, our findings are negative for all three questions, indicating that the assumptions in the proposed argument do not hold for syntactic structures. Therefore, it is needed to reconsider the relationship between the power law and hierarchical structures.
In addition, we perform the same analyses on trees generated by the PCFG that approximates syntactic structures, to compare them with natural languages. These analyses highlight the differences between PCFGs and natural languages. Furthermore, their statistical properties are also different from those assumed in the argument by Lin and Tegmark (2017). This means that the PCFG does not converge to its asymptotic behavior within the scale of a single sentence. Consequently, it is difficult to apply the proposed argument not only to single sentences, but also to other types of short sequences such as child languages and non-human signals.
To investigate if the argument holds beyond the sentence level, future research will need to formalize the description of hierarchical structures of entire texts, such as discourse. Otherwise, we need to seek for an alternative argument that account for the power-law behavior.
2 Background
2.1 Power-law Decay of Correlation
[…]
In contrast, the power-law decay, $C\sim r^{-\alpha}$, is slower than the exponential one for any correlation length $\xi$, meaning that the correlation is significant for any long distances. In other words, the scale of correlation diverges. Moreover, power-law functions are scale-invariant. For the scale transformation $r^{\prime}=kr$, we can rescale the correlation by $C^{\prime}=k^{-\alpha}C$ so that the relation remains the same, $C^{\prime}\sim r^{\prime-\alpha}$. This property also implies the divergence of the scale.
Therefore, the power-law decay observed in various corpora means that the scale diverges in natural languages. Intuitively, a single element change can significantly influence the overall structure of a text.
2.2 Power Law and Hierarchical Structures
In the study of human languages and animal signals, the power-law decay of correlation has been interpreted as evidence of underlying hierarchical structures. This interpretation is based on the work of Lin and Tegmark (2017), which shows that the decay of correlation follows a power law in sequences generated by PCFGs. PCFGs generate tree structures from the root to leaves iteratively, where the state of a parent node determines the states of its children. The leaves correspond to a sequence, while the whole tree is the underlying hierarchical structure. PCFGs are a probabilistic extension of context-free grammars (CFGs; Chomsky, 1956). CFGs were originally introduced to describe the nested structures, such as the if…then construction. Due to their simplicity and ability to capture fundamental aspects of hierarchical syntactic structures, PCFGs have been used to model syntactic structures (Charniak, 1997). Beyond linguistics, PCFGs have also been applied to represent hierarchical structures in diverse domains (Ellis et al., 2015; Worth and Stepney, 2005; Gilbert and Conklin, 2007; Knudsen and Hein, 1999; Harlow et al., 2012; Li, 1989b, 1991; Tano et al., 2020; Worth and Stepney, 2021).
Figure 1: Sequential and structural distances. (a) If trees are balanced, the sequential distance grows exponentially with the structural distance, $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$. (b) If trees are strongly biased, the growth is slower. For example, the relation is linear, $r_{\text{seq}}\sim r_{\text{str}}$.
We give an overview of the discussion by Lin and Tegmark (2017). To understand their discussion, it is useful to distinguish between sequential and structural distances, which we refer to as $r_{\text{seq}}$ and $r_{\text{str}}$, respectively. The former is defined along the sequence. On the other hand, the latter is the distance in the hierarchical structure behind the sequence, as shown in Figure 1.
Lin and Tegmark (2017) have proven that the correlation asymptotically decays according to an exponential function of the structural distance, $C\sim\exp(-\lambda r_{\text{str}})$. They also assumed that generated trees are balanced, i.e., they have neither left- nor right-branching bias, as in Figure 1 (a). Letting the depth of the two nodes from the common ancestor be $d$, the structural distance is roughly $r_{\text{str}}\sim 2d$, and the sequential distance is asymptotically exponential with $d$, $r_{\text{seq}}\sim q^{d}$, where $q$ is the typical number of children of each node. Therefore, the sequential distance grows exponentially with the structural distance, $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$, where $\mu=(\ln q)/2$. Combining $C\sim\exp(-\lambda r_{\text{str}})$ and $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$ leads to the power-law decay of correlation in sequences, $C\sim r_{\text{seq}}^{-\alpha}$, with $\alpha=\lambda/\mu$.
[…]
It is widely accepted among linguists that the syntactic structure of a sentence is hierarchical. Moreover, formal descriptions of these structures have been well established. Therefore, this study aims to examine the validity of the proposed argument, focusing on syntactic structures.
Meanwhile, the proposed argument is not limited to syntactic structures in principle. It may also apply to other types of structures, such as semantic and discourse ones. For example, entire texts composed of many sentences intuitively appear to have hierarchical structures defined by paragraphs, sections, chapters, and so on. This hierarchical organization may explain why the decay of correlation follows a power law at scales beyond the sentence level, extending across entire texts. The structures may also be more abstract. Previous research has observed the power-law decay in sequences without syntax, such as child languages and animal signals, arguing that these also have underlying hierarchical structures (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
However, the existence and the formalization of these types of hierarchical structures remain unclear. Therefore, we are currently unable to assess whether the proposed argument accounts for the power law at scales beyond the sentence level or in early-stage or non-human languages.
2.3 Rethinking Previous Research
The argument by Lin and Tegmark (2017) raises several important questions that must be addressed to determine whether it can apply to syntactic structures. First, do hierarchical structures in natural languages actually exhibit the behaviors assumed in their argument? More specifically:
  1. i)
    Does the correlation in hierarchical structures decay exponentially with the structural distance?
  2. ii)
    Are the hierarchical structures sufficiently balanced such that the sequential distance grows exponentially with the structural distance?
Particularly, the latter appears to contradict the well-known fact that syntactic structures typically exhibit a left- or right-branching bias (Hawkins, 1990; Dryer, 1992) (for instance, see Figure 2). In such cases, the growth is slower than exponential. For example, Figure 1 (b) illustrates a case where the growth is linear, $r_{\text{seq}}\sim r_{\text{str}}$.
Figure 2: Example of a parse tree from the BLLIP corpus, preprocessed as described in Section 3.1. This tree has a strong right-branching bias.
Moreover, it is unlikely that PCFGs model syntactic structures sufficiently well. Linguistically, some languages cannot be described by CFGs (Shieber, 1985; Culy, 1985; Senuma and Aizawa, 2024). In addition, empirically, naive parsers based on PCFGs are insufficient for parsing (Johnson et al., 2006; O’Donnell et al., 2009). If syntactic structures deviate from PCFGs significantly, the discussion by Lin and Tegmark (2017) based on PCFGs should not be applied to them. This leads to the last question:
  1. iii)
    Do PCFGs describe the statistical properties of hierarchical structures sufficiently?
In this study, we perform statistical analyses on parse trees to address these questions.
3 Methods
3.1 Datasets and Preprocessing
We adopt the Brown Laboratory for Linguistic Information Processing 1987-89 Corpus Release 1 (BLLIP; Charniak et al., 2000), a large-scale English treebank that provides Penn Treebank (Marcus et al., 1993)-style syntactic annotations for three years’ worth of Wall Street Journal articles. This treebank contains approximately 30 million words and one million sentences in total. In advance, we remove the phonological null elements such as traces of movement and ellipsis to ensure the alignment with the WikiText-103 corpus (Merity et al., 2017) analyzed in Appendix D, which does not include traces.
The number of child nodes differs across nodes in the dateset, making it difficult to align nodes across different parse trees. To address this issue, we binarize the trees using the Tree.chomsky_normal_form method from the NLTK library (Bird and Loper, 2004). We then apply the Tree.collapse_unary method from the same library to remove the unary branches produced by the former method. In addition, because a larger number of possible node states requires more data for reliable statistical estimation, we replace phrasal and part-of-speech (POS) tags with a smaller set of categories. Specifically, we refer to the Universal Dependencies (de Marneffe et al., 2021), which contains fewer types of tags than the Penn Treebank-style corpus, and group tags that tend to have similar types of dependents and heads. We present the eleven categories used in our analyses in Appendix A. An example of a preprocessed parse tree is shown in Figure 2.
3.2 Mutual Information
As a metric for correlations, we use MI as in Lin and Tegmark (2017). The MI between two tags is defined by
$I=\sum_{x_{0},x_{1}}P(x_{0},x_{1})\ln\frac{P(x_{0},x_{1})}{P(x_{0})P(x_{1})},$ (1)
where $P(x_{0})$ and $P(x_{1})$ are the marginal probabilities of the first and second tags, and $P(x_{0},x_{1})$ is their joint probability. A naive estimate of $I$ can be obtained by substituting the empirical distributions $\hat{P}(x_{0})$, $\hat{P}(x_{1})$, and $\hat{P}(x_{0},x_{1})$ into Eq. (1). However, this estimation is biased positively due to finite samples (Li, 1990). To reduce this bias, we decompose the MI into
$I$ $=S_{0}+S_{1}-S_{01},$ (2)
where $S_{0}$, $S_{1}$, and $S_{01}$ are the Shannon entropy, defined by $S_{0}=-\sum_{x_{0}}P(x_{0})\ln P(x_{0})$, $S_{1}=-\sum_{x_{1}}P(x_{1})\ln P(x_{1})$, and $S_{01}=-\sum_{x_{0},x_{1}}P(x_{0},x_{1})\ln P(x_{0},x_{1})$. Then we estimate the entropy $S_{0}$ using the method proposed by Grassberger (2003):
$\hat{S}_{0}=\psi(N)-\frac{1}{N}\sum_{x_{0}}n_{x_{0}}\psi(n_{x_{0}}),$ (3)
where $\psi$ is the digamma function, $n_{x_{0}}$ is the number of samples such that the former tag is $x_{0}$, and $N=\sum_{x_{0}}n_{x_{0}}$ is the total number of samples. $S_{1}$ and $S_{01}$ are estimated in a similar manner. By computing $\hat{S}_{0}$, $\hat{S}_{1}$, and $\hat{S}_{01}$ from the empirical distributions and using Eqs. (2) and (3), we can estimate the MI with a smaller bias.
To evaluate the sufficiency of the sampling, we compute the MI estimate while varying the number of samples, $N_{\text{data}}$. If the estimate converges as $N_{\text{data}}$ increases, this suggests that sampling bias is negligible.
To estimate MI in sequences, we sample $N_{\text{data}}$ pairs of POS nodes separated by a sequential distance $r_{\text{seq}}$, from the dataset. For MI in structures, we sample $N_{\text{data}}$ pairs of nodes at a fixed structural distance $r_{\text{str}}$, where the nodes are either POS or phrasal. We treat ordered pairs as distinct to avoid arbitrarily selecting one of the two. For example, we count pairs (AUX, PP) and (PP, AUX), corresponding to were and than, in Figure 2 as distinct to each other if both pairs are sampled.
3.3 Growth of Sequential Distance
To address question (ii), we investigate how the sequential distance depends on the structural distance. Specifically, for each $r_{\text{seq}}$ and $r_{\text{str}}$, we count the number of POS-tagged node pairs over the whole treebank whose sequential and structural distances are $r_{\text{seq}}$ and $r_{\text{str}}$, respectively. We also compute the average of sequential distance for each structural distance $r_{\text{str}}$.
3.4 Context-free Independence Breaking
To examine the validity of PCFGs as models of syntactic structures, it is necessary to quantify the extent to which structures deviate from PCFGs. We employ a metric for this quantification, called context-free independence breaking (CFIB; Nakaishi and Hukushima, 2024).
The essential property of PCFGs, which we refer to as the context-free independence, is that the distribution of states of each node depends only on its parent. Consequently, given the states of two nodes, their children are mutually independent.
Therefore, we can quantify the deviation from PCFGs by the MI between two pairs under the condition that the states of their parents are fixed, as presented in Figure 3. Specifically, provided that the states of two nodes are $x_{0}$ and $x_{1}$, the MI between their children is defined as
$\begin{split}J_{x_{0}x_{1}}=\sum_{y_{0},z_{0},z_{0},z_{1}}P(y_{0},z_{0},y_{1},z_{1}|x_{0},x_{1})\\ \times\ln\frac{P(y_{0},z_{0},y_{1},z_{1}|x_{0},x_{1})}{P(y_{0},z_{0}|x_{0},x_{1})P(y_{1},z_{1}|x_{0},x_{1})}.\end{split}$ (4)
Here, $P(y_{0},z_{0}|x_{0},x_{1})$ denotes the conditional probability that the left and right children of the node with label $x_{1}$ are $y_{0}$ and $z_{0}$, respectively. Similarly, $P(y_{1},z_{1}|x_{0},x_{1})$ refers to those for the node with label $x_{1}$. $P(y_{0},z_{0},y_{1},z_{1}|x_{0},x_{1})$ is the conditional joint probability of both pairs of children. By construction, this metric is non-negative for any tree structures and zero for PCFGs.
Figure 3: CFIB $J$ is the MI between the children of two nodes whose states are fixed.
For the estimation of the CFIB, we sample $N_{\text{data}}$ pairs of nodes that are tagged as $x_{0}$ and $x_{1}$, with the structural distance of $r_{\text{str}}$. Then, we compute the empirical distributions of tags on the children of the nodes. From these distributions, we estimate $J$ in the same procedure explained in the previous section. We also distinguish between pairs and their permutations, as in the estimation of MI.
3.5 PCFG Approximation
We also construct a PCFG that approximates the distribution of parse trees. This allows for a more detailed comparison between syntactic structures and PCFGs. The PCFG is constructed using the maximum likelihood estimation, which is equivalent to simply counting the frequency of each production rule in the dataset. Using the PCFG derived from the preprocessed dataset from BLLIP, we generate trees $1.6\times 10^{7}$ times, with up to $1\times 10^{2}$ iterations per generation. This process resulted in $13,256,547$ terminated trees. We then apply the same analyses to these generated trees as to parse trees.
3.6 Fitting
We test whether the decay of correlation follows an exponential or power-law form by fitting the data with an exponential function, $A\exp(-\lambda r)$, and a power-law function, $Br^{-\alpha}$. To perform the fitting, we apply the least squares method after log-scaling the MI and CFIB values, because fitting with a linear scale puts too much weight on data points at shorter distances. We also fit the growth of sequential distance using an exponential function, $C\exp(\mu r)$, and a power-law function, $Dr^{\beta}$, without log-scaling the values. For both fitting, we use the lmfit Python package Newville et al. (2014). In the following results, we present the fitted parameters $A$, $\lambda$, and others along with the reduced chi-square statistic, $\chi^{2}_{\nu}$.
4 Results
4.1 Mutual Information in Sequences
First, we confirm that the correlation in sequences decays according to a power law. Figure 4 shows the MI $I_{\text{POS}}$ between POS tags as a function of the sequential distance $r_{\text{seq}}$ for different data sizes $N_{\text{data}}$. The estimated values converge as $N_{\text{data}}$ increases, indicating that the estimation bias is sufficiently small. In addition, the decay is better fitted by a power-law function than by an exponential one. This result indicates that the correlation between POS tags within a sentence follows a power law, consistent with previous research. Similar results are also observed for the WikiText corpus, see Appendix D.
Figure 4: MI $I_{\text{POS}}$ between POS tags as a function of the sequential distance $r_{\text{seq}}$. The fitted exponential and power-law decays for $N_{\text{data}}=2.56\times 10^{6}$ are also presented.
4.2 Mutual Information in Structures
In this subsection, we address question (i), whether the exponential decay of correlation with the structural distance holds in syntactic structures. The estimated MI between POS or phrasal tags as a function of the structural distance is shown in Figure 5. As in the previous subsection, the convergence of the estimated values with increasing $N_{\text{data}}$ indicates that the bias is negligible. Interestingly, model fitting reveals that the correlation decays according to a power law rather than an exponential form with respect to the structural distance. This result indicates that syntactic structures do not meet the condition necessary for the previously proposed argument to hold. We have also confirmed that similar results are obtained for the unbinarized case and the WikiText corpus, as detailed in Appendices C and D.
Figure 5: MI $I_{\text{tag}}$ between POS or phrasal tags as a function of the structural distance $r_{\text{str}}$. The fitted exponential and power-law decays for $N_{\text{data}}=2.56\times 10^{6}$ are also presented.
4.3 Growth of Sequential Distance
The second question is (ii): Does the sequential distance grow exponentially with increasing structural distance? To examine this, we count the number of POS tag pairs such that the structural and sequential distances are $r_{\text{str}}$ and $r_{\text{seq}}$, respectively. The results are shown in Figure 6. We also compute the averages sequential distance for each structural distance, represented by black circles, and fit the results using both exponential and power-law functions. The fitting reveals that the growth follows a power law. However, the trend becomes less clear for larger structural distances $r_{\text{str}}\gtrsim 50$ because of data sparsity. Notably, the power law exponent $\beta=0.87$ is smaller than $1$, indicating that the growth is sublinear, i.e., slower than linear. As demonstrated in Figure 1 (b), this slow growth quantitatively reflects a storing branching bias in syntactic structures. These findings clearly show that the second condition does not hold for syntactic structures. We have also observed similar results for the unbinarized case and for the WikiText corpus in Appendices C and D.
Figure 6: Number of POS tag pairs such that the structural and sequential distances are $r_{\text{str}}$ and $r_{\text{seq}}$, respectively. Black circles represent the average sequential distance for each structural distance.
4.4 Context-free Independence Breaking
Question (iii) asks whether PCFGs are valid models for the hierarchical structures of natural languages. In this section, we measure the CFIB $J$ as the metric of deviations from PCFGs for syntactic structures. In Figure 7, we present the results for fixed nodes being $(x_{0},x_{1})=(\text{NP},\text{NP})$, the most frequent pair in the corpus. The convergence of the estimates with increasing $N_{\text{data}}$ confirms that the estimation bias is negligible. Moreover, the CFIB values are clearly positive, meaning that syntactic structures do not have the essential property of PCFGs. Furthermore, the metric decays according to a power law with respect to structural distance, suggesting that the context-free independence is significantly broken even at large distances. These findings lead to answering question (iii) in the negative. As shown in Appendix B, similar behaviors are observed for other frequent node pairs $(x_{0},x_{1})=(\text{NP},\text{VP})$ and $(x_{0},x_{1})=(\text{VP},\text{VP})$. In addition, Appendix D shows that the decay of $J$ also follows a power law for the WikiText corpus.
Figure 7: CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$ as a function of the structural distance $r_{\text{str}}$. The fitted exponential and power-law decays for $N_{\text{data}}=2.56\times 10^{6}$ are also presented.
4.5 PCFG approximation
To deepen our understanding of the results above in comparison with PCFGs, we apply the same analyses to the PCFG constructed to approximate the BLLIP corpus. Figure 8 summarizes the results. The most striking difference between the parse trees and the PCFG appears in the CFIB metric, shown in Figure 8 (d). The estimated value of this metric does not saturate but continues to decrease even for large data sizes. This behavior is consistent with the fact that CFIB is identically zero for PCFGs by definition.
For the other statistical quantities, the differences are more moderate. As shown in Figure 8 (a), the MI in sequences decays according to a power law at short sequential distances, $r_{\text{seq}}\lesssim 10$, while for large sequential distances, $r_{\text{seq}}\gtrsim 10$, it increases substantially, in contrast to the trend in parse trees. The MI in structures, as shown in Figure 8 (b), exhibits a power-law decay as in parse trees. In Figure 8 (c), the average sequential distance grows more slowly than exponentially and even than linearly, indicating that the generated trees are strongly biased. While this behavior is qualitatively similar to that in syntactic structures, the estimated exponent $\beta=0.45$ significantly differs from that for the parse trees, where $\beta=0.87$.
Figure 8: Statistical properties of trees generated by the PCFG obtained from the maximum likelihood estimation of syntactic structures. (a) MI $I_{\text{POS}}$ between POS tags. (b) MI $I_{\text{tag}}$ between POS or phrasal tags. (c) Distribution and average of sequential distances. (d) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$. Fitting was performed for $N_{\text{data}}=8\times 10^{6}$ in (a) and (b), and for $r_{\text{str}}\leq 10$ in (a).
Although these findings may suggest that the PCFG partially captures the statistical properties of parse trees, the mechanism underlying the power-law decay of correlation in sequences clearly differs from the one proposed in Lin and Tegmark (2017), in which the decay of MI in structures and the growth of sequential distance are exponential. Importantly, our finding that the MI in structures decays according to a power law in the PCFG suggests that the distance within a sentence is too short for the MI to converge to its asymptotic exponential decay.
This indicates that the proposed argument is difficult to apply to short sequences such as single sentences. In such cases, it becomes impossible to distinguish whether the process underlying the sequences is Markov-like or PCFG-like. This limitation also arises when analyzing child languages and animal languages, because available datasets do not contain sequences of sufficient length. Indeed, previous studies have analyzed sequences shorter than $\sim 10^{2}$ phones, syllables, words, or action elements (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
5 Discussion and Conclusion
Lin and Tegmark (2017) attributed the power-law decay of correlation in natural languages to underlying hierarchical structures, which has inspired numerous studies on both human languages and animal signals. However, their explanation raises several important questions, as discussed in Section 2.3. To address these, the present study analyzes the statistical properties of hierarchical syntactic structures in natural languages, to test the questions essential for the proposed argument.
The analyses answer all the questions negatively: (i) The correlation in syntactic structures decays according to a power law with respect to the structural distance; (ii) The growth of the sequential distance is slower than linear with the structural distance, due to the strong branching bias in syntactic structures; (iii) The statistical properties of syntactic structures differ from those of PCFGs, particularly in the CFIB. These results reveal that the argument by Lin and Tegmark (2017) cannot account for the mechanism underlying the power law at the sentence level.
We also analyze the PCFG that approximates the parse trees. The behavior of CFIB in the PCFG clearly differs from that in the parse trees. Furthermore, this analysis demonstrates that the asymptotic behavior of the PCFG does not emerge for short sequences. This result highlights the difficulty of applying the proposed argument not only to single sentences in adult human languages but also to child languages and animal signals.
While further investigation is needed for other parsing schemes and corpora, we expect qualitatively similar results in more general settings. Indeed, we have confirmed similar results for unbinarized parse trees and for the WikiText corpus (Merity et al., 2017), see Appendices C and D for details. Whether these behaviors appear in other languages remains an open question. A key challenge is the limited availability of treebanks for languages other than English.
Another important open question is whether the argument by Lin and Tegmark (2017) applies at scales longer than sentences, that is, at the level of entire texts, where the statistical behaviors of PCFGs may be sufficiently close to their asymptotic behaviors. To extend our analysis to the level, a necessary first step is to formalize how to describe the hierarchical structures of entire texts and develop datasets annotated with such information.
Since the discussion in previous research does not account for the power law in natural languages, at least at the sentence level, alternative explanations are also needed. One promising direction is critical phenomena in statistical physics. These refer to the phenomena that statistical quantities exhibit power-law behaviors near phase transition point, where statistical properties of a system change qualitatively. Indeed, as the temperature parameter of large language models varies, generated texts exhibit critical phenomena, sharing statistical properties with natural languages, including the power-law decay of correlation (Nakaishi et al., 2024).
Finally, more detailed characterizations of the statistical quantities are necessary. This study used simple fitting functions to distinguish between exponential and power-law behaviors. However, the observed statistical quantities exhibit more complex patterns. A closer examination of these patterns will be essential for understanding syntactic structures through statistical analysis.
Acknowledgments
This work was supported by JSPS KAKENHI Grant Nos. 23H01095, 23KJ0622, and 24H00087; JST, PRESTO Grant No. JPMJPR21C2; and the World-Leading Innovative Graduate Study Program for Advanced Basic Science Course at the University of Tokyo.
Appendix A Tag Reduction for Parse Trees
To reduce the number of node states, we grouped tags with similar functions into the same category. The eleven categories used in our analyses are shown in Table 1.
Tags Part-of-Speech Tags and Phrase Labels
NP NP NN NNP NNS CD PRP TMP POS QP NNPS EX NX FW
VP VP VBD VB VBN VBG VBZ VBP TO
S S S1 SBAR SINV FRAG SQ SBARQ RRC
PP IN PP RP PRT
DT DT PRP\$ PDT
JJ JJ ADJP JJR JJS NAC ADJ
Others , . ” “ \$ : PRN -RRB- -LRB- # X INTJ UH SYM LS LST HYPH NFP
AUX AUX MD AUXG
RB RB ADVP RBR RBS
CC CC UCP CONJP
WH WHNP WDT WP WHADVP WRB WHPP WP\$ WHADJP NML
Table 1: Eleven tags used in this study and the corresponding part-of-speech tags and phrase labels.
Appendix B Context-free Independence Breaking
[…]
Appendix C Unbinarized Case
As discussed in Section 3.1, we binarized syntactic trees during preprocessing. To verify that this binarization does not significantly affect our results, we analyze the statistical properties of the unbinarized trees in this section. The MI in sequences is unaffected by binarization. In addition, the dataset is too limited to reliably compute the CFIB for unbinarized trees. Therefore, we focus on the MI in structures and the growth of sequential distance. The results, shown in Figure 10, exhibit the power-law decay of the MI and the slow growth of the sequential distance, which are qualitatively consistent with those observed in the binarized case, although, as expected, the structural distances are smaller. These suggest that the statistical behavior of syntactic structures is robust to binarization.
Figure 9: CFIB $J$ for (a) $(x_{0},x_{1})=(\text{NP},\text{VP})$ and (b) $(x_{0},x_{1})=(\text{VP},\text{VP})$. Fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$ in (a) and $N_{\text{data}}=6.4\times 10^{5}$ in (b).
Figure 10: Results for unbinarized parse trees. (a) MI $I_{\text{tag}}$ between POS or phrasal tags. Fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$. (b) Distribution and average of sequential distances.
Figure 11: Results for WikiText. (a) MI $I_{\text{POS}}$ between POS tags. (b) MI $I_{\text{tag}}$ between POS or phrasal tags. (c) Distribution and average of sequential distances. (d) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$. Fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$ in (a), (b), and (d).
Appendix D WikiText
In the main text, we analyzed the BLLIP corpus. To address whether similar behaviors are observed in other corpora, we perform the same analyses for the WikiText-103 corpus (wikitext-103-raw-v1 split) (Merity et al., 2017), which is a collection of English Wikipedia articles that preserves the original Wikipedia markup and contains over 100 million tokens and approximately 3.9 million sentences. We implement a rule-based detokenizer and segment sentences with a spaCy sentencizer (Honnibal et al., 2020). Then, we extract sentences up to 40 words (resulting in approximately 3.4 million sentences) and apply the Berkeley Neural Parser (benepar_en3 model) (Kitaev and Klein, 2018; Kitaev et al., 2019) to them to obtain Penn Treebank-style constituency trees. Finally, we binarize the trees and reduce the number of node states using the same method applied to the BLLIP corpus.
Figure 11 shows that the results are qualitatively similar to those for BLLIP: the MI decays according to a power-law function in both sequences and structures; the growth of sequential distance is slower than linear; and the CFIB also decays following a power law. These suggest that the observed behaviors are universal, and that the assumption in Lin and Tegmark (2017) does not hold for syntactic structures across corpora.
However, we note a slight difference in the behavior of MI in both sequences and structures: it increases at large distances, $r_{\text{seq}}\gtrsim 30$ and $r_{\text{str}}\gtrsim 20$. This may be due to structural differences in unusually long sentences or potential parsing errors in such cases. We are unable to determine the cause of this effect, as the dataset contains too few long sentences.