### v2

# Rethinking the Relationship between the Power Law and Hierarchical Structures

Abstract
Statistical analysis of corpora provides an approach to quantitatively investigate natural languages. This approach has revealed that several power laws consistently emerge across different corpora and languages, suggesting universal mechanisms underlying languages. Particularly, the power-law decay of correlation has been interpreted as evidence for underlying hierarchical structures in syntax, semantics, and discourse. This perspective has also been extended to child speeches and animal signals. However, the argument supporting this interpretation has not been empirically tested in natural languages. To address this problem, the present study examines the validity of the argument for syntactic structures. Specifically, we test whether the statistical properties of parse trees align with the assumptions in the argument. Using English and Japanese corpora, we analyze the mutual information, deviations from probabilistic context-free grammars (PCFGs), and other properties in natural language parse trees, as well as in the PCFG that approximates these parse trees. Our results indicate that the assumptions do not hold for syntactic structures and that it is difficult to apply the proposed argument to child speeches and animal signals, highlighting the need to reconsider the relationship between the power law and hierarchical structures.
11footnotetext: This work was conducted while KN and $K$ were at the University of Tokyo.
1 Introduction
Natural languages exhibit characteristic statistical properties. Quantifying these properties through corpus-based analyses is crucial for understanding natural languages, providing insights into the mechanisms that give rise to the properties. Previous research has extensively explored this approach, uncovering various power laws, such as Zipf’s law (Zipf, 1949) and Heaps’ law (Heaps, 1978), that consistently appear across different corpora and languages. These findings suggest the existence of universal mechanisms underlying these power laws.
This study focuses on the power-law decay of correlation. Previous studies have measured correlations between phones, characters, words, and other elements using metrics such as correlation functions and mutual information (MI), across various corpora. These studies have consistently found that the correlation decays according to a power-law function of the distance between elements (Li, 1989a; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023; Nakaishi et al., 2024).
The power-law decay is qualitatively slower than the exponential decay, which is typically observed when interactions between elements are local, as in Markov chains. Therefore, the power-law decay of correlation in natural languages indicates that distant elements remain strongly correlated. This observation is consistent with linguistic intuition. For example, syntactic dependencies can, in principle, span arbitrarily long distances through repeated center-embedding, and a word introduced early in a context may still influence interpretation much later.
Moreover, the power-law decay of correlation has been interpreted as evidence for the linguistic view that languages are organized hierarchically in syntax, semantics, and discourse (Alvarez-Lacalle et al., 2006; Altmann et al., 2012; Lin and Tegmark, 2017). This interpretation has also been extended beyond speeches and written texts by human adults. Previous research has observed similar power-law behaviors in child speeches and in animal signals, arguing that these may also possess underlying hierarchical structures (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
This interpretation is justified by Lin and Tegmark (2017). They showed that the decay of correlation follows a power law in sequences generated by probabilistic context-free grammars (PCFGs; Jelinek et al., 1992), which are a simple model for generating hierarchical tree structures. To be more precise, they proved that in PCFGs, the correlation decays exponentially in structures. In addition, they assumed that the sequential distance grows exponentially with the structural distance. These two give rise to the power-law decay of correlation in sequences.
However, several assumptions underlying this argument seem implausible in light of well-established linguistic facts. Yet, previous studies have not tested the validity of these assumptions quantitatively using empirical data.
To address this, we examine whether these assumptions hold for hierarchical structures in natural language syntax, i.e., parse trees. Syntactic structures have been well formalized, resulting in the availability of large-scale treebanks. This enables us to perform statistical analyses on treebanks to test these assumptions.
Specifically, we formalize three key questions essential for examining the validity of the assumptions: (i) Does the correlation decay exponentially in syntactic structures?; (ii) Does the sequential distance grow exponentially with the structural distance?; (iii) To what extent do the statistical properties of syntactic structures deviate from those of PCFGs? To answer these questions, we analyze English and Japanese treebanks.
As a result, our findings are negative for all three questions, indicating that the assumptions in the proposed argument do not hold for syntactic structures in natural languages. Therefore, it is necessary to reconsider the relationship between the two established facts, the power-law decay of correlation and the hierarchical syntactic structures in natural languages.
In addition, we perform the same analyses on trees generated by the PCFG that approximates parse trees, to highlight differences between natural language syntactic structures and those generated by the PCFG. These analyses further show that the MI in structures generated by the PCFG does not exhibit exponential decay when the length of generated sequences is comparable to that of single sentences. Although Lin and Tegmark (2017) proved that the decay converges to an exponential form at sufficiently large distances, which their argument relies on, the decay follows a non-exponential form at distances typically observed in natural language parse trees. The parse trees are too small for this asymptotic exponential behavior to emerge. Consequently, it is difficult to apply their argument not only to single sentences by adults but also child speeches and non-human signals, in which the sequential distance is typically as short as in single sentences.
While the present study focuses on the hierarchical syntactic structures of individual sentences, the structures of entire texts consisting of multiple sentences, such as discourse structures, also appear to be hierarchical. The argument by Lin and Tegmark (2017), if applied to these structures, might account for the power-law decay of correlation in natural languages.
It would therefore be fruitful to investigate the relationship between the power-law behavior and hierarchical structures at the discourse level. Although previous studies have proposed formal frameworks for discourse annotation (Miltsakaki et al., 2004; Zeldes et al., 2025), the size of existing corpora remain too small to examine the validity of the argument. Developing sufficiently large-scale discourse corpora would be an important next step for future research.
Another direction is to explore alternative mechanisms that could give rise to power-law behavior in natural languages.
2 Background
2.1 Power-law Decay of Correlation
Previous research has shown that the correlation in natural languages decays according to a power-law function of distance. This phenomenon has been observed across various corpora and languages (Li, 1989a; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023; Nakaishi et al., 2024).
[…]
In contrast, the power-law decay is expressed as $C\sim r^{-\alpha}$. Although the right-hand side resembles a polynomial form, the exponent $\alpha$ is a real number rather than a natural number. This decay is slower than the exponential one for any correlation length $\xi$, meaning that the correlation is significant for any long distances. In other words, the scale of correlation diverges.
Moreover, power-law functions are scale-invariant. For the scale transformation $r^{\prime}=kr$, we can rescale the correlation by $C^{\prime}=k^{-\alpha}C$ so that the relation remains the same, $C^{\prime}\sim r^{\prime-\alpha}$. This property also implies the divergence of the scale.
Therefore, the power-law decay of correlation observed in various corpora means that the scale diverges in natural languages. Intuitively, this implies that a local change can propagate and affect the global structure of a text.
2.2 Power Law and Hierarchical Structures
In studies of human languages and animal signals, this power-law behavior has been interpreted as evidence of underlying hierarchical structures. This interpretation is based on the work of Lin and Tegmark (2017), which shows that the decay of correlation follows a power law in sequences generated by PCFGs.
PCFGs generate tree structures in a top-down manner, from the root to the leaves iteratively, where the category of a parent node determines the categories of its children. The leaves correspond to a surface sequence, while the whole tree represents the underlying hierarchical structure.
PCFGs are a probabilistic extension of context-free grammars (Chomsky, 1956), which were originally introduced to describe the nested structures, such as the if…then construction. PCFGs have also been used to model syntactic structures (Charniak, 1997), due to their simplicity and ability to capture fundamental aspects of hierarchical syntactic structures.
We provide an overview of the discussion in Lin and Tegmark (2017). To understand their discussion, it is useful to distinguish between sequential and structural distances, which we refer to as $r_{\text{seq}}$ and $r_{\text{str}}$, respectively. The former is defined along the sequence. The latter is the distance in the hierarchical structure behind the sequence, as shown in Figure 1.
Figure 1: Sequential and structural distances. (a) If trees are balanced, the sequential distance grows exponentially with the structural distance, $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$. (b) If trees are strongly biased, the growth is slower. For example, the relation is linear, $r_{\text{seq}}\sim r_{\text{str}}$.
Lin and Tegmark (2017) have proven that in PCFGs, the correlation decays exponentially with the structural distance, $C\sim\exp(-\lambda r_{\text{str}})$, at sufficiently large distances. Note that the decay may follow non-exponential forms at shorter distances. They also assumed that generated trees are balanced, i.e., the trees have neither left- nor right-branching bias, as in Figure 1 (a). Letting the depth of the two nodes from the common ancestor be $d$, the structural distance is roughly $r_{\text{str}}\sim 2d$, and the sequential distance is asymptotically exponential with $d$, $r_{\text{seq}}\sim q^{d}$, where $q$ is the typical number of children of each node. Therefore, the sequential distance grows exponentially with the structural distance, $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$, where $\mu=(\ln q)/2$. Combining $C\sim\exp(-\lambda r_{\text{str}})$ and $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$ leads to the power-law decay of correlation in sequences, $C\sim r_{\text{seq}}^{-\alpha}$, with $\alpha=\lambda/\mu$.
[…]
The present study aims to examine the validity of this proposed argument. Throughout this study, we focus on syntactic structures. It is widely accepted in linguistics that these structures are hierarchical, with well-established formal descriptions. Using large-scale datasets based on this formalism, we can test if the proposed argument holds for syntactic structures.
It is worth noting that, in principle, PCFGs are models not only for syntactic structures but also for more general hierarchical structures. Indeed, PCFGs have been applied to represent hierarchical structures in diverse domains (Ellis et al., 2015; Worth and Stepney, 2005; Gilbert and Conklin, 2007; Knudsen and Hein, 1999; Harlow et al., 2012; Li, 1989b, 1991; Tano et al., 2020; Worth and Stepney, 2021).
From this perspective, the argument by Lin and Tegmark (2017) may also apply to other types of structures, such as semantic and discourse ones. For example, entire texts composed of many sentences intuitively appear to have hierarchical structures defined by paragraphs, sections, chapters, and so on. This hierarchical organization may explain why the decay of correlation follows a power law beyond individual sentences, extending across entire texts.
However, it is difficult to empirically test whether the argument holds for multi-sentence texts. Although previous studies have proposed formal frameworks for annotating discourse structures (Miltsakaki et al., 2004; Zeldes et al., 2025), existing discourse corpora are not yet large enough to reliably estimate the statistical quantities introduced in Section 3, which we use to examine the validity of the assumptions in Lin and Tegmark (2017). Reliable estimation of these quantities requires substantial amounts of data.
2.3 Rethinking Previous Research
To test whether the argument by Lin and Tegmark (2017) holds for syntactic structures in natural languages, several important questions must be addressed.
First, do syntactic structures in natural languages actually exhibit the behaviors assumed in their argument? More specifically:
  1. i)
    Does the correlation in syntactic structures decay exponentially with the structural distance?
  2. ii)
    Are syntactic structures sufficiently balanced such that the sequential distance grows exponentially with the structural distance?
Figure 2: Example of a parse tree from BLLIP, preprocessed as described in Section 3.1. This tree has a strong right-branching bias.
Particularly, the second one appears to contradict the well-known fact that syntactic structures typically exhibit a left- or right-branching bias (Hawkins, 1990; Dryer, 1992) (for instance, see Figure 2). In such cases, the growth can be slower than exponential. For example, Figure 1 (b) illustrates a case where the growth is linear, $r_{\text{seq}}\sim r_{\text{str}}$.
In addition, Lin and Tegmark (2017) demonstrate their argument using PCFGs as models of hierarchical structures in natural languages. However, the distribution that PCFGs can represent is highly limited. In this model, the distribution of a subtree depends only on its root, a property we refer to as the context-free independence. Previous studies have suggested that relaxing this property is necessary for modeling syntactic structures (Collins, 2003; Johnson et al., 2006; O’Donnell et al., 2009). Therefore, the last question is:
  1. iii)
    Do the statistical properties of syntactic structures deviate significantly from those of PCFGs?
These three questions remain unresolved. First, although extensive research has measured the correlation in texts and speeches, the correlation in the underlying structures have not been quantified. Therefore, the answer to question (i) remains elusive. Second, syntactic structures exhibit branching bias. Previous studies have not quantitatively examined how this bias affects the relationship between structural and sequential distances, which is crucial for addressing question (ii). Finally, it is well known that PCFGs cannot adequately model natural language syntax without being augmented to relax the context-free independence. However, to answer question (iii) clearly, we need to systematically quantify the differences in the statistical properties between natural language syntax and PCFGs.
In this study, we perform statistical analyses on parse trees to address these three questions, one by one.
3 Methods
3.1 Datasets and Preprocessing
We primarily use the Brown Laboratory for Linguistic Information Processing 1987-89 Corpus Release 1 (BLLIP; Charniak et al., 2000). Its large scale enables highly precise estimation of the statistical quantities.
To verify the robustness of our results, we also conduct the same analyses on two additional corpora. The first is WikiText (Merity et al., 2017), an English corpus from a different domain than BLLIP. The second is the NINJAL Parsed Corpus of Modern Japanese (NPCMJ; National Institute for Japanese Language and Linguistics, 2016), a treebank of Japanese, which is typologically distinct from English.
BLLIP
BLLIP (Charniak et al., 2000) is a large-scale English treebank that provides Penn Treebank (Marcus et al., 1993)-style syntactic annotations for three years’ worth of Wall Street Journal articles. This treebank contains approximately 30 million words and one million sentences in total. We remove the phonological null elements, such as traces of movement and ellipsis, to ensure alignment with WikiText-103 (Merity et al., 2017), which does not include traces.
The number of child nodes differs across nodes in the dataset, making it difficult to align nodes across different parse trees. To address this issue, we binarize the trees using the Tree.chomsky_normal_form method implemented in the NLTK library (Bird and Loper, 2004). We then apply the Tree.collapse_unary method from the same library to remove the unary branches produced by the former method.
In addition, because a larger number of possible categories requires more data for reliable statistical estimation, we replace part-of-speech (POS) and phrasal tags with a smaller set of categories. Specifically, we refer to the Universal Dependencies (de Marneffe et al., 2021), which contains fewer types of tags than the Penn Treebank-style corpus, and group tags that tend to have similar types of dependents and heads. The eleven categories used in this study are listed in Appendix A. We apply the same categories to both POS and phrasal tags, as the argument by Lin and Tegmark (2017) treats leaves and internal nodes in the same manner. An example of a preprocessed parse tree is shown in Figure 2.
The statistical quantities analyzed in the present study are defined for the distribution of trees and estimated from the empirical dataset. Because parse trees depend on the chosen annotation scheme, these quantities also depend on the scheme. Nevertheless, we expect that the main results, such as whether the decay is exponential or power-law, will remain unchanged as long as the annotation scheme is reasonable.
To verify this, we also perform the same analyses under two alternative settings: trees without binarization and the phrasal-tag-only setting, in which we distinguish phrasal tags from POS tags and focus only on the former. Appendix B shows that the results for both settings are qualitatively similar to those presented in the main text, that is, the answers to questions (i), (ii), and (iii) are negative.
WikiText
To examine whether our conclusions hold beyond BLLIP, we perform the same analyses on WikiText-103 (wikitext-103-raw-v1 split) (Merity et al., 2017), which consists of English Wikipedia articles that preserves the original Wikipedia markup and contains over 100 million tokens and approximately 3.9 million sentences. We implement a rule-based detokenizer and segment sentences with a spaCy sentencizer (Honnibal et al., 2020). Then, we extract sentences up to 40 words (resulting in approximately 3.4 million sentences) and apply the Berkeley Neural Parser (benepar_en3 model) (Kitaev and Klein, 2018; Kitaev et al., 2019) to them to obtain Penn Treebank-style constituency trees. Finally, we binarize the trees and reduce the number of categories using the same method applied to BLLIP.
NPCMJ
We also analyze the Japanese manually-annotated treebank, NPCMJ (National Institute for Japanese Language and Linguistics, 2016). It contains more than 1.4 million tokens and approximately 90,000 sentences. We first omit all of the phonologically null tokens (null pronouns, traces, etc.) and extract sentences of up to 40 words, resulting in approximately 82,000 sentences. Then we binarize all parse trees and systematically rewrite the POS and phrasal tags using a rule-based mapping scheme listed in Appendix A. Since Japanese is a head-final langauge, all trees are binarized in a left-branching direction, in contrast to the right-branching binarization applied to English.
3.2 Mutual Information
As a metric for correlations, we use the mutual information (MI) as in Lin and Tegmark (2017). The MI between two tags is defined by
$I=\sum_{x_{0},x_{1}}P(x_{0},x_{1})\ln\frac{P(x_{0},x_{1})}{P(x_{0})P(x_{1})},$ (1)
where $P(x_{0})$ and $P(x_{1})$ are the marginal probabilities of the first and second tags, and $P(x_{0},x_{1})$ is their joint probability. The MI can be decomposed into
$I$ $=S_{0}+S_{1}-S_{01},$ (2)
where $S_{0}$, $S_{1}$, and $S_{01}$ are the Shannon entropy, defined by $S_{0}=-\sum_{x_{0}}P(x_{0})\ln P(x_{0})$, $S_{1}=-\sum_{x_{1}}P(x_{1})\ln P(x_{1})$, and $S_{01}=-\sum_{x_{0},x_{1}}P(x_{0},x_{1})\ln P(x_{0},x_{1})$. By estimating $S_{0}$, $S_{1}$, and $S_{01}$, we can estimate the MI.
A naive estimate can be obtained by substituting the empirical distributions $\hat{P}(x_{0})$, $\hat{P}(x_{1})$, and $\hat{P}(x_{0},x_{1})$ into $P(x_{0})$, $P(x_{1})$, and $P(x_{0},x_{1})$ in the definitions of entropy. In this case, the estimates of entropy and MI are biased due to finite samples (Arora et al., 2022; Li, 1990).
To reduce this bias, we employ the more sophisticated estimator proposed by Grassberger (2003). In this method, the entropy $S_{0}$ is estimated by
$\hat{S}_{0}=\psi(N)-\frac{1}{N}\sum_{x_{0}}n_{x_{0}}\psi(n_{x_{0}}),$ (3)
where $\psi$ is the digamma function, $n_{x_{0}}$ is the number of samples such that the former tag is $x_{0}$, and $N=\sum_{x_{0}}n_{x_{0}}$ is the total number of samples. $S_{1}$ and $S_{01}$ are estimated in a similar manner. Grassberger (2003) has proven that the bias of this estimator rapidly converges to zero as the number of samples, $N$, increases. Therefore, if the estimated value saturates with respect to $N$, the finite-sample bias should be negligible.
To estimate the MI in sequences, we sample $N_{\text{data}}$ pairs of POS nodes separated by a sequential distance $r_{\text{seq}}$, from the dataset. For the MI in structures, we sample $N_{\text{data}}$ pairs of nodes at a fixed structural distance $r_{\text{str}}$, where the nodes are either POS or phrasal. We treat ordered pairs as distinct to avoid arbitrarily selecting one of the two. For example, we count pairs (AUX, PP) and (PP, AUX), corresponding to were and than, in Figure 2 as distinct to each other if both pairs are sampled.
In the following analyses, we compute the MI estimates while varying the number of samples, $N_{\text{data}}$. In our results, the estimates saturate as $N_{\text{data}}$ increases, indicating that the bias is sufficiently small.
To further support our claim, we also report the results obtained with six alternative entropy estimators Madow (1948); Miller (1955); Zahl (1977); Horvitz and Thompson (1952); Chao and Shen (2003); Wolpert and Wolf (1995); Nemenman et al. (2001) in Appendix C. The results are consistent with those reported in the main text, except for the estimators that were previously reported to exhibit larger errors (Arora et al., 2022).
3.3 Growth of Sequential Distance
To address question (ii), we investigate how the sequential distance depends on the structural distance. Specifically, for each $r_{\text{seq}}$ and $r_{\text{str}}$, we count the number of POS-tagged node pairs over the whole treebank whose sequential and structural distances are $r_{\text{seq}}$ and $r_{\text{str}}$, respectively. We also compute the average of sequential distance for each structural distance $r_{\text{str}}$.
3.4 Context-free Independence Breaking
To address question (iii), it is necessary to quantify the extent to which syntactic structures deviate from PCFGs. We employ a metric for this quantification, called context-free independence breaking (CFIB; Nakaishi and Hukushima, 2024).
The essential property of PCFGs, which we refer to as the context-free independence, is that the distribution of categories of each node depends only on its parent. Consequently, given the categories of two nodes, their children are mutually independent.
Figure 3: CFIB $J$ is the MI between the children of two nodes whose categories are fixed.
Therefore, we can quantify the deviation of the distribution of trees from PCFGs by the MI between two pairs under the condition that the categories of their parents are fixed, as presented in Figure 3. Specifically, provided that the categories of two nodes are $x_{0}$ and $x_{1}$, the CFIB is defined as
$\begin{split}J_{x_{0}x_{1}}=\sum_{y_{0},z_{0},z_{0},z_{1}}P(y_{0},z_{0},y_{1},z_{1}|x_{0},x_{1})\\ \times\ln\frac{P(y_{0},z_{0},y_{1},z_{1}|x_{0},x_{1})}{P(y_{0},z_{0}|x_{0},x_{1})P(y_{1},z_{1}|x_{0},x_{1})}.\end{split}$ (4)
Here, $P(y_{0},z_{0}|x_{0},x_{1})$ denotes the conditional probability that the left and right children of the node with label $x_{1}$ are $y_{0}$ and $z_{0}$, respectively. Similarly, $P(y_{1},z_{1}|x_{0},x_{1})$ refers to those for the node with label $x_{1}$. $P(y_{0},z_{0},y_{1},z_{1}|x_{0},x_{1})$ is the conditional joint probability of both pairs of children. By definition, the CFIB is non-negative for any tree structures and zero for trees generated by PCFGs.
The CFIB is defined for the distribution of trees, as is the MI. We can estimate this metric from the dataset of trees, similarly to the estimation of the MI. We sample $N_{\text{data}}$ pairs of nodes that are tagged as $x_{0}$ and $x_{1}$, with the structural distance of $r_{\text{str}}$. Then, we compute the empirical distributions of tags on the children of the nodes. From these distributions, we estimate $J$ in the same procedure explained in the previous section. We also distinguish between pairs and their permutations, as in the estimation of MI.
We verify the accuracy of the estimates for the CFIB by varying $N_{\text{data}}$, as we did for the MI. Appendix C also presents the results obtained with alternative estimators, further supporting the validity of our estimates.
3.5 PCFG Approximation
We also construct a PCFG that approximates the distribution of natural language parse trees. This allows for a more detailed comparison between syntactic structures and PCFGs. The PCFG is constructed using the maximum likelihood estimation, which is equivalent to simply counting the frequency of each production rule in the dataset. Using the PCFG derived from the preprocessed dataset from BLLIP, we generate trees $1.6\times 10^{7}$ times, with up to $1\times 10^{2}$ iterations per generation.This procedure yields approximately 13 million terminated trees. We then apply the same analyses to these generated trees as to parse trees.
3.6 Fitting
We test whether the decay of correlation follows an exponential or power-law form by fitting the data with an exponential function, $A\exp(-\lambda r)$, and a power-law function, $Br^{-\alpha}$. To perform the fitting, we apply the least squares method after log-scaling the MI and CFIB values, because fitting with a linear scale puts too much weight on data points at shorter distances. We also fit the growth of sequential distance using an exponential function, $C\exp(\mu r)$, and a power-law function, $Dr^{\beta}$, without log-scaling the values. For both fitting, we use the lmfit Python package Newville et al. (2014). In the following results, we present the fitted parameters $A$, $\lambda$, and others along with the reduced chi-square statistic, $\chi^{2}_{\nu}$.
3.7 Synthetic Data
To demonstrate that our method can reliably estimate MI and distinguish between exponential and power-law decays, we apply it to mathematical models in which MI can be computed exactly. We generate random variables $X$ and $Y$ with joint probabilities defined as $P(X=0,Y=0)=P(X=1,Y=1)=(1+\delta)/4$ and $P(X=0,Y=1)=P(X=1,Y=0)=(1-\delta)/4$. The MI between $X$ and $Y$ is $I=\frac{1+\delta}{2}\ln(1+\delta)+\frac{1-\delta}{2}\ln(1-\delta)$, which is $\delta^{2}/2+\mathcal{O}(\delta^{3})$ as $\delta\to 0$. If $\delta=\exp(-\lambda r/2)$, the MI decays asymptotically as $I\sim\exp(-\lambda r)$. If $\delta=r^{-\alpha/2}$, the decay follows a power law, $I\sim r^{-\alpha}$. We refer to these as the exponential and power-law models, respectively.
Figure 4: Estimation and fitting of the MI for (a) the exponential model with $\lambda=0.1$ and (b) the power-law model with $\alpha=2$. Fitting was performed for $N_{\text{data}}=8\times 10^{6}$.
We sample $X$ and $Y$ from the exponential model with $\lambda=0.1$ and from the power-law model with $\alpha=2$. We then estimate the MI and fit the resulting values. The results are shown in Figure 4. In both cases, the estimates converge to the exact values as the sample size $N_{\text{data}}$ increases. In addition, $\chi^{2}_{\nu}$ is significantly smaller when the fitting function matches the model. Our method successfully discriminates between the exponential and power-law models.
4 Results
4.1 Mutual Information in Sequences
First, we confirm that the correlation in sequences decays according to a power law. Figure 5 shows the MI $I_{\text{POS}}$ between POS tags as a function of the sequential distance $r_{\text{seq}}$ for different data sizes $N_{\text{data}}$. The estimated values converge as $N_{\text{data}}$ increases, indicating that the estimation bias is sufficiently small.
The decay of $I_{\text{POS}}$ is better fitted by a power-law function ($\chi^{2}_{\nu}=2.6\times 10^{-1}$) than by an exponential one ($\chi^{2}_{\nu}=9.8\times 10^{-1}$). This result indicates that the correlation between POS tags follows a power law, consistent with previous research.
Figure 5: MI $I_{\text{POS}}$ between POS tags as a function of the sequential distance $r_{\text{seq}}$. The fitted exponential and power-law decays for $N_{\text{data}}=2.56\times 10^{6}$ are also presented.
4.2 Mutual Information in Structures
We now address question (i): whether the exponential decay of correlation with the structural distance holds in natural language syntactic structures. The estimated MI between POS or phrasal tags as a function of the structural distance is shown in Figure 6. As in the previous section, the convergence of the estimated values with increasing $N_{\text{data}}$ indicates that the bias is negligible.
Model fitting reveals that the correlation decays according to a power law ($\chi^{2}_{\nu}=1.1\times 10^{-1}$) rather than an exponential form ($\chi^{2}_{\nu}=1.3$) with respect to the structural distance. This result indicates that syntactic structures do not meet the condition necessary for the previously proposed argument to hold.
Figure 6: MI $I_{\text{tag}}$ between POS or phrasal tags as a function of the structural distance $r_{\text{str}}$. The fitted exponential and power-law decays for $N_{\text{data}}=2.56\times 10^{6}$ are also presented.
4.3 Growth of Sequential Distance
Question (ii) asks whether the sequential distance grows exponentially with increasing structural distance. To examine this, we count the number of POS tag pairs such that the structural and sequential distances are $r_{\text{str}}$ and $r_{\text{seq}}$, respectively. The results are shown in Figure 7. We also compute the average sequential distance for each structural distance, represented by black circles, and fit the results using both exponential and power-law functions.
The fitting reveals that the growth follows a power law ($\chi^{2}_{\nu}=1.4\times 10$) rather than an exponential form ($\chi^{2}_{\nu}=1.4\times 10^{21}$), although the trend becomes less clear for larger structural distances $r_{\text{str}}\gtrsim 50$ because of data sparsity. Notably, the power law exponent $\beta=0.72$ is smaller than $1$, indicating that the growth is slower than linear. As demonstrated in Figure 1 (b), this slow growth quantitatively reflects a strong branching bias in syntactic structures. These findings clearly show that the second condition does not hold for syntactic structures.
Figure 7: Number of POS tag pairs such that the structural and sequential distances are $r_{\text{str}}$ and $r_{\text{seq}}$, respectively. Black circles represent the average sequential distance for each structural distance.
4.4 Context-free Independence Breaking
Question (iii) concerns whether the statistical properties of syntactic structures deviate significantly from those of PCFGs. We measure the CFIB $J$ as a metric of this deviation. In Figure 8, we present the results for fixed nodes being $(x_{0},x_{1})=(\text{NP},\text{NP})$, the most frequent pair in the corpus. The convergence of the estimates with increasing $N_{\text{data}}$ confirms that the estimation bias is negligible.
The CFIB values are clearly positive, meaning that syntactic structures do not have the essential property of PCFGs. Furthermore, the metric decays according to a power law ($\chi^{2}_{\nu}=8.1\times 10^{-2}$) rather than an exponential form ($\chi^{2}_{\nu}=7.1\times 10^{-1}$) with respect to structural distance, suggesting that the context-free independence is significantly broken even at large distances. These results indicate that the answer to question (iii) is negative.
This finding is also consistent with previous studies that developed PCFG-based parsers achieving high performance (Collins, 2003; Johnson et al., 2006; O’Donnell et al., 2009). These studies augmented PCFGs to relax the context-free independence implicitly or explicitly.
Figure 8: CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$ as a function of the structural distance $r_{\text{str}}$. The fitted exponential and power-law decays for $N_{\text{data}}=2.56\times 10^{6}$ are also presented.
As shown in Appendix D, similar behaviors are observed for other frequent node pairs $(x_{0},x_{1})=(\text{NP},\text{VP})$ and $(x_{0},x_{1})=(\text{VP},\text{VP})$, suggesting that the observed deviation from PCFGs is not specific to a particular syntactic category pair.
4.5 Other Corpora
We examined whether the three assumptions in the argument of Lin and Tegmark (2017) hold for BLLIP. The results show that all of these assumptions are violated. The argument cannot account for the statistical properties of parse trees in this corpus. To test whether this conclusion is specific to BLLIP, we conduct the same analyses on WikiText and NPCMJ.
Figure 9: Statistical properties of syntactic structures in WikiText and NPCMJ. (a1–4) Results for WikiText: (a1) $I_{\text{POS}}$ as a function of $r_{\text{seq}}$; (a2) $I_{\text{tag}}$ as a function of $r_{\text{str}}$; (a3) relationship between $r_{\text{seq}}$ and $r_{\text{str}}$; and (a4) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$ as a function of $r_{\text{str}}$. Fitting was performed with $N_{\text{data}}=2.56\times 10^{6}$. (b1–b4) Corresponding results for NPCMJ. Fitting was performed with $N_{\text{data}}=6.4\times 10^{5}$ for $I_{\text{POS}}$ and $I_{\text{tag}}$, and with $N_{\text{data}}=4.0\times 10^{4}$ for $J$.
Figure 9 summarizes the results. For both corpora, the statistical properties are similar to those observed for BLLIP: the MI decays according to a power law in both sequences and structures; the growth of sequential distance is slower than linear; and the CFIB also decays following a power law. These behaviors are supported by the chi-square statistics, which are smaller for power-law fitting than for exponential fitting across all statistical quantities.
The results for these corpora also exhibit slight differences from those for BLLIP. $I_{\text{POS}}$ and $I_{\text{tag}}$ for WikiText, as well as $I_{\text{POS}}$ for NPCMJ, increase at large distances. This effect may be attributable to structural differences in unusually long sentences or to potential parsing errors in such cases. However, we cannot determine the significance of this behavior because the datasets contain too few long sentences.
Nonetheless, the results clearly support that the assumptions proposed by Lin and Tegmark (2017) are violated not only for BLLIP, but also for another English corpus, WikiText, as well as a Japanese treebank, NPCMJ. These results imply the robustness of our conclusion, given that Japanese is typologically distinct from English, exhibiting head-final, left-branching syntactic structures.
4.6 PCFG Approximation
Figure 10: Statistical properties of trees generated by the PCFG obtained from the maximum likelihood estimation of syntactic structures. (a) MI $I_{\text{POS}}$ between POS tags. (b) MI $I_{\text{tag}}$ between POS or phrasal tags. (c) Distribution and average of sequential distances. (d) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$. Fitting was performed for $N_{\text{data}}=8\times 10^{6}$ in (a) and (b), and for $r_{\text{str}}\leq 10$ in (a).
To further elucidate the difference between syntactic structures and PCFGs, we apply the same analyses to the PCFG constructed using maximum likelihood estimation on BLLIP. The results are summarized in Figure 10.
Mutual Information in Sequences
Figure 10 (a) shows the MI in sequences. At short sequential distances, $r_{\text{seq}}\lesssim 10$, the MI decays according to a power law, although the decay rate does not clearly match that observed in parse trees. At larger distances, $r_{\text{seq}}\gtrsim 10$, the MI increases substantially, in contrast to the trend in parse trees.
Mutual Information in Structures
The MI in structures, shown in Figure 10 (b), exhibits a power-law decay similar to that in syntactic structures, although the decay rate differs. This result does not contradict the properties of PCFGs proven by Lin and Tegmark (2017). While the MI in structures generated by PCFGs asymptotically follows an exponential decay at sufficiently large distances, the MI can exhibit different behavior at shorter distances. The non-exponential behavior of $I_{\text{tag}}$ suggests that for individual sentences, the structural distances are too short for the asymptotic behavior to emerge.
If the generation process continued for more steps, producing sufficiently large trees, the MI in Figure 10 (b) would converge to an exponential decay at large structural distances. However, such a setup cannot be justified because the generated trees would be much larger than typical parse trees. Even if we assume this setup, the MI in sequences would decay more rapidly than a power law as long as the trees are sufficiently biased so that the sequential distances grow more slowly than exponentially. The mechanism of the power-law decay of correlation observed in natural languages remains unexplained.
The non-exponential behavior of $I_{\text{tag}}$ in the PCFG has another important implication. As discussed in Section 2.2, the argument by Lin and Tegmark (2017) can, in principle, be extended to hierarchical structures beyond syntax. Following this perspective, several studies have applied the argument to domains other than texts and speeches by human adults. These studies reported the power-law decay of correlation in child speeches and animal signals, interpreting it as evidence for the existence of underlying hierarchical structures in these domains (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
However, the argument by Lin and Tegmark (2017) relies on the asymptotic behavior of PCFGs, which does not appear when the observed sequences are too short. In child speeches and animal signals, the sequences are typically shorter than sentences produced by human adults. In such cases, it is difficult to interpret the power-law decay of correlation as evidence for hierarchical structures.
Growth of Sequential Distance
In Figure 10 (c), the average sequential distance grows more slowly than exponentially and even than linearly, indicating that the generated trees are strongly biased. While this behavior is qualitatively similar to that in parse trees, the estimated exponent $\beta=0.41$ significantly differs from that for the parse trees, where $\beta=0.72$.
Context-free Independence Breaking
The most striking difference between parse trees and trees generated by the PCFG appears in the CFIB metric. The true value of this metric is zero for PCFGs by definition. If the estimates converge to this true value with increasing data size, the corresponding values in the log-scale plot continue to decrease, approaching -$\infty$, since the logarithm of zero is -$\infty$.
In Figure 10 (d), the estimates of CFIB is much smaller than those for syntactic structures in Figure 8. Furthermore, the estimates continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, which is the expected behavior. This result is in contrast to that for syntactic structures in natural languages.
5 Discussion and Conclusion
In natural languages, the correlation decays as a power law with respect to the sequential distance, as reported in numerous previous studies. At the same time, linguists generally agree that natural languages exhibit hierarchical syntactic structures. Although both phenomena are well established, their relationship remains elusive.
The present study examines this relationship, specifically the argument proposed by Lin and Tegmark (2017). They attributed the power-law decay of correlation in natural languages to underlying hierarchical structures, which has since inspired numerous studies on both human languages and animal signals. However, their argument relies on several assumptions that are questionable from a linguistic perspective. To examine the validity of the assumptions for syntactic structures, we test the three key questions through statistical analyses of parse trees in natural languages.
The analyses yield negative answers to all three questions: (i) The correlation in syntactic structures decays according to a power law with respect to the structural distance; (ii) The growth of the sequential distance is slower than linear with the structural distance, due to the strong branching bias in syntactic structures; (iii) The statistical properties of syntactic structures differ from those of PCFGs, particularly in the CFIB. These results reveal that the relationship between the power-law decay of correlation and syntactic structures differs from that proposed by Lin and Tegmark (2017).
We obtained similar results under various settings: for the English treebank BLLIP with several different schemes, for another English corpus, WikiText, and for the Japanese treebank NPCMJ. The consistency of the results across different schemes, domains, and languages suggests the robustness of our conclusion.
We also analyze the PCFG obtained from the maximum likelihood estimation of BLLIP. The statistical properties of the PCFG clearly differ from those in parse trees. Furthermore, this analysis demonstrates that the exponential decay of $I_{\text{tag}}$ in the PCFG does not emerge within individual sentences. Although Lin and Tegmark (2017) proved that $I_{\text{tag}}$ decays exponentially at sufficiently large structural distances, the structural distances in typical parse trees are too short for this behavior to appear. This result highlights the difficulty of applying the proposed argument not only to sentences produced by human adults but also to child speeches and animal signals, in which the sequences are shorter than sentences by human adults.
In addition, the statistical differences between natural language parse trees and trees generated by PCFGs may have implications for theoretical research on large language models. Some studies in this field analyze models trained on synthetic data, in which PCFGs are sometimes used to generate the data (Allen-Zhu and Li, 2023; Cagnetta and Wyart, 2024). However, the model’s ability to learn PCFGs will not account for its ability to learn syntactic structures of natural languages, although some studies employ PCFGs as models of more abstract structures than syntax.
Currently, we cannot rule out the possibility that the MI in natural language syntactic structures follows an exponential decay at structural distances much larger than those observed in this study. However, even if this is the case, such large parse trees would be too rare to have a significant influence on the statistical properties of the corpus as a whole. Moreover, as long as syntactic structures are biased, the decay of MI in sequences should be faster than a power law, because the sequential distance grows more slowly than exponentially. Therefore, statistical analyses of larger parse trees are unlikely to reveal the mechanism underlying the power-law decay of correlation in natural languages.
A more promising direction is to consider structures larger than individual sentences, namely, the hierarchical organization of entire texts composed of multiple sentences. The argument by Lin and Tegmark (2017) may account for the power-law decay of correlation observed across entire texts rather than within single sentences. As a first step in this direction, it will be necessary to develop sufficiently large-scale corpora annotated with discourse information to estimate statistical quantities such as $I_{\text{POS}}$, $I_{\text{tag}}$, and $J$.
Another direction is to explore approaches different from that of Lin and Tegmark (2017). One alternative is critical phenomena in statistical physics. These refer to the phenomena that statistical quantities exhibit power-law behaviors near phase transition points, at which statistical properties of a system change qualitatively. Indeed, as the temperature parameter of large language models varies, generated texts exhibit critical phenomena, sharing statistical properties with natural languages, including the power-law decay of correlation (Nakaishi et al., 2024).
Acknowledgments
We thank Taiga Ishii, Shinnosuke Isono, Sho Yokoi, and Keisuke Sakaguchi for useful discussions and suggestions. This work was supported by JSPS KAKENHI Grant Nos. 23H01095, 23KJ0622, 24H00087, and 25K24434; JST, PRESTO Grant No. JPMJPR21C2; and the World-Leading Innovative Graduate Study Program for Advanced Basic Science Course at the University of Tokyo.
Appendix A Tag Reduction
To reduce the number of categories, we grouped tags with similar functions into the same category. The eleven categories used for BLLIP and WikiText and the seven categories for NPCMJ are shown in Tables 1 and 2, respectively.
Tags Part-of-Speech Tags and Phrase Labels
NP NP NN NNP NNS CD PRP TMP POS QP NNPS EX NX FW
VP VP VBD VB VBN VBG VBZ VBP TO
S S S1 SBAR SINV FRAG SQ SBARQ RRC
$P$ IN PP RP PRT
DT DT PRP\$ PDT
$J$ JJ ADJP JJR JJS NAC ADJ
Others , . ” “ \$ : PRN -RRB- -LRB- # X INTJ UH SYM LS LST HYPH NFP
AUX AUX MD AUXG
RB RB ADVP RBR RBS
$C$ CC UCP CONJP
WH WHNP WDT WP WHADVP WRB WHPP WP\$ WHADJP NML
Table 1: Eleven tags used for BLLIP and WikiText and the corresponding part-of-speech tags and phrase labels.
Tags Part-of-Speech Tags and Phrase Labels
NP NP NUMCLP PRN NML PNLP NUMCLPSYM N NPR NUM CL ADJN PRO ADJI D Q FN WPRO PNL PRN WNUM WD FW
IP IP VB AX AXD VB0 VB2 PASS PASS2 MD
$P$ PP P
ADVP ADVP ADV NEG WADV
CONJP CONJP CONJ
$P$ PP P
CP CP FRAG INTJP FS INTJ
Others LST META LS PU PUL PUR SYM PUQ COMMENT
Table 2: Seven tags used for NPCMJ and the corresponding part-of-speech tags and phrase labels.
Appendix B Other Schemes
In the main text, we analyzed parse trees after applying two preprocessing steps. First, we binarized the parse trees. Second, we reduced the number of tag types by grouping similar tags, where we use the same reduced categories for both POS and phrasal tags.
Although the statistical quantities depend on the specific preprocessing scheme, we expect that these details do not affect our conclusions, such as whether the MI decays according to a power law or an exponential function, and whether the CFIB is significantly positive. To examine this expectation, we perform additional analyses under two alternative preprocessing schemes: the unbinarized setting and the phrasal-tag-only setting, in which POS and phrasal tags are distinguished and only the latter are analyzed.
Figure 11: (a1, 2) Results for the unbinarized setting: (a1) MI $I_{\text{tag}}$ between POS or phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) distribution and average of sequential distances. (b1, 2) Results for the phrasal-tag-only setting: (b1) MI $I_{\text{tag}}$ between phrasal tags, where fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$; (a2) CFIB $J$, where fitting was pwerformed for $N_{\text{data}}=6.4\times 10^{5}$.
Figures 11 (a1) and (a2) show the results for the unbinarized setting. Because binarization does not affect the MI in sequences and the dataset is too limited to reliably estimate the CFIB for unbinarized trees, we present only the MI in structures and the sequential distance as a function of the structural distance. For both quantities, the results are qualitatively similar to those for the binarized setting in the main text: the decay of $I_{\text{tag}}$ follows a power-law function rather than an exponential function, and the growth of $r_{\text{seq}}$ is slower than exponential.
In the phrasal-tag-only setting, the MI between POS tags and the sequential distance are not defined. We therefore estimate the MI between phrasal tags and the CFIB, presented in Figures 11 (b1) and (b2), respectively. Both quantities exhibit power-law decay, consistent with the results in the main text.
Across both alternative preprocessing schemes, our main conclusion holds: the assumptions in the argument by Lin and Tegmark (2017) are not satisfied. These results indicate that our findings are robust to the details of preprocessing.
Appendix C Alternative Estimators
In the main text, we computed the MI using the estimator of Grassberger (2003) (GR), whose bias decreases rapidly. The estimated values saturate as $N_{\text{data}}$ increases, except for the CFIB in the PCFG. The estimates of the CFIB in the PCFG continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, which is reasonable because the true value is zero. These results show that we have properly estimated the statistical quantities.
To further support the reliability of our estimates, we compute $I_{\text{POS}}$, $I_{\text{tag}}$, and $J$ using six alternative estimators proposed by Chao and Shen (2003) (CS), Horvitz and Thompson (1952) (HT), Miller (1955) and Madow (1948) (MM), Nemenman et al. (2001) (NSB), Wolpert and Wolf (1995) (WW) and Zahl (1977) (ZA). For these computations, we use the Python package entropy-estimation (Arora et al., 2022).
Figure 12 presents the results for BLLIP computed with the six alternative estimators, together with those obtained using the GR estimator, which are the same as in the main text. The estimates agree with ours in most cases. Figures 12 (b) and (c) show that the WW estimates for $I_{\text{tag}}$ and $J$ at large distances exhibit clear bias, producing some negative estimates even though the true MI and CFIB are always non-negative. The CS and HT estimates for $J$ also exhibit similar behaviors. These results are consistent with Arora et al. (2022), which reports that these three estimators yield larger errors than the others when the sample size is large. From these results, we can confirm that our estimates in the main text are reliable.
Figure 12: Results for BLLIP obtained with different estimators. (a) MI $I_{\text{POS}}$ between POS tags. (b) MI $I_{\text{tag}}$ between POS or phrasal tags. (c) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$. In all cases, $N_{\text{data}}=2.56\times 10^{6}$.
Figure 13: Results for the PCFG obtained with different estimators. (a) MI $I_{\text{POS}}$ between POS tags. (b) MI $I_{\text{tag}}$ between POS or phrasal tags. In both (a) and (b), $N_{\text{data}}=8\times 10^{6}$. (c-e) CFIB $J$ for $(x_{0},x_{1})=(\text{NP},\text{NP})$ obtained with (c) $M$, and (d) NSB, and (e) ZA.
Figure 13 shows the results for the same quantities for the PCFG obtained from the maximum likelihood estimation of BLLIP. The estimates of $I_{\text{POS}}$ and $I_{\text{tag}}$, shown in Figures 13 (a) and (b), are consistent across all estimators. This again supports the reliability of the results in the main text.
We also show the estimates of the CFIB $J$ obtained with the MM, NSB, and ZA estimators in Figures 13 (c), (d), and (e), respectively. The MM and ZA estimates continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, similar to the GR estimates in the main text. This means that these two estimates properly converge to zero, the true value of the CFIB in PCFGs. The convergence of the NSB estimates is rather unclear. We do not present results with CS, HT, and WW, because most of their estimates are negative and cannot be plotted on a log scale.
In summary, MM, ZA, and GR, which is used in the main text, perform well for all statistical quantities. NSB generally performs well, yet its convergence to the true value is unclear for the CFIB in the PCFG. The other estimators, CS, HT, and WW, sometimes exhibit clear bias, implying that they should be avoided when the sample size is large.
Appendix D Context-free Independence Breaking
In Section 4.4, we presented the results of CFIB for BLLIP when the two nodes are fixed as $(x_{0},x_{1})=(\text{NP},\text{NP})$. In this section, we present the results with $(x_{0},x_{1})=(\text{NP},\text{VP})$ and $(\text{VP},\text{VP})$ to examine how the metric depends on the fixed nodes.
From the results shown in Figure 14, we can observe a power-law decay similar to that with $(x_{0},x_{1})=(\text{NP},\text{NP})$, although the power law exponents may be different. This implies that the CFIB in syntactic structures follows a power-law decay, irrespective of the fixed nodes.
Figure 14: CFIB $J$ for BLLIP with (a) $(x_{0},x_{1})=(\text{NP},\text{VP})$ and (b) $(x_{0},x_{1})=(\text{VP},\text{VP})$. Fitting was performed for $N_{\text{data}}=2.56\times 10^{6}$ in (a) and $N_{\text{data}}=6.4\times 10^{5}$ in (b).